An Angel Investor has been asked to invest $200,000 into a new firm. The Angel Investor expects to realize an annual rate of return of 25% over a 4 year period. The estimate value of the firm in four years is believed to be $6 million. Based on these facts what is the dollar value of the return the Angel Investor expect to receive and what percentage of ownership does this amount equate to for the Angel Investor? Please explain how you arrived at the numbers.
Answer #1

The dollar value of return after 4 years =FV-Investment =200000*(1+25%)^4-200000=288281.25

Value of investment after 4 years =200000*(1+25%)^4 =488281.25

Percentage of ownership =Value of Investment/Estimated Value of firm =488281.25/6000000=8.14%